OpenStudy (anonymous):

how to find tan 5pi/4 on the unit circle

OpenStudy (jdoe0001):

use eyes

OpenStudy (anonymous):

I tried that and it didn't work

OpenStudy (jdoe0001):

\(\bf \cfrac{5}{4}\implies \cfrac{1}{4}+\cfrac{4}{4}\qquad {\color{blue}{ thus}}\qquad \cfrac{5\pi}{4}\implies \cfrac{\pi}{4}+\cfrac{4\pi}{4}\implies \cfrac{\pi}{4}+\pi\)

OpenStudy (jdoe0001):

see it?

OpenStudy (anonymous):

I see but where would it be placed on the unit circle

OpenStudy (jdoe0001):